# AI/ML Based Detection and Categorization of Covert Communication in IPv6 Network

### v1

@@ 5.3.4 Accuracy
6 Results
6.1 Comparative Performance Analysis of Deployed Models
The performance of various machine learning models in the covert communication detection task was rigorously evaluated using precision, recall, $F_{1}$ score, and accuracy and can be observed in Figure 5 and Figure 6. These metrics serve as fundamental indicators of each model’s performance and are depicted in the accompanying bar charts for both binary and multiclass classification scenarios.
6.1.1 Binary Classification
With excellent performance on all criteria, the Random Forest (RF), Extreme Gradient Boosting (XGB), and Light Gradient Boosting (LGB) models proved their capacity to discriminate between the two groups. The RF and XGB models outperformed the others in terms of accuracy and $F_{1}$ scores, demonstrating their resilience in binary classification tasks.
The Graph Convolutional Network (GCN), on the other hand, had much lower results, indicating that it had trouble adjusting to binary classification. Because non-Euclidean data structures are less important in smaller binary classification tasks, GCN’s reliance on them may be the cause of this underperformance. With lower scores than other neural network models, the Bidirectional LSTM (LSTM) likewise demonstrated modest performance, suggesting that it may be improved in binary classification settings.
6.1.2 Multiclass Classification
In the more challenging multiclass classification task, RF and XGB models maintained their high performance, further emphasizing their adaptability to complex scenarios involving multiple classes. The Light Gradient Boosting (LGB) model outperformed all other models, achieving the highest accuracy and $F_{1}$ scores, which highlights its exceptional capability in handling the intricacies of multiclass classification.
Interestingly, the Bidirectional LSTM (LSTM) emerged as a particularly effective model for multiclass classification tasks, demonstrating a remarkable ability to balance precision and recall. This performance can likely be attributed to the LSTM’s inherent capacity to capture sequential dependencies and patterns in the data. In particular, the LSTM appears well-suited for identifying specific covert communication classes, such as Hop Limit encoded packets, where detecting the sequence of changes in hop limit values is crucial for accurate classification. By leveraging its sequence modeling capabilities, the LSTM was able to detect subtle temporal and sequential patterns that are often missed by other architectures.
Table 3: Model Performance: Accuracy and $F_{1}$ Score for Binary and Multiclass Classification
@@ 6.1.2 Multiclass Classification
However, the GCN model’s performance declined significantly, indicating that it was not very suitable for multiclass classification in this situation. The model’s natural optimization for tasks involving graph-structured data, as opposed to feature-rich datasets typical of covert communication detection, may be the cause of this.
The Deep Neural Network (DNN), Convolutional Neural Network (CNN), and Recurrent Neural Network (RNN) showed balanced and competitive performance in multiclass classification, with CNN slightly leading among neural network models. However, the standout performance of LSTM underscores its strength in understanding and leveraging temporal sequences to differentiate between complex covert communication types.
Overall, traditional tree-based models such as RF, XGB, and especially LGB demonstrated exceptional performance across both binary and multiclass classification tasks, reaffirming their robustness for covert communication detection. Neural network-based models also performed competitively, though GCN’s architectural design constraints led to suboptimal results. Table 3 summarizes the accuracy and $F_{1}$ scores for both tasks.
6.2 Generative AI Assisted Refinement
In this work, we refined Python scripts related to covert communication detection using the GPT-4-turbo model as a generative AI agent expert in code generation. Utilizing generative AI, the classifier models underwent iterative refinement, with a binary classifier for detection and a multi-class classifier for identifying the type of covert communication. The main objective of the generative AI agents was to iteratively refine the Python scripts until performance exceeded a predetermined threshold. Figure 7 illustrates the process.
Figure 7: Prompt Engineering Example for Script Refinement
The system prompt includes detailed information about the dataset, such as column properties, specific IPv6 fields vulnerable to covert communication, and other contextual information that may impact the detection process. The user prompt contains the path to the dataset, classification results (provided as data frames), the Python scripts needing refinement, and the classification accuracy threshold. Generative AI makes systematic improvements by iteratively reviewing, adjusting, and honing the classifiers until they reach or surpass the accuracy threshold. This process of refining leads to improved detection and classification accuracy in cybersecurity applications by effectively automating model selection and script improvements utilizing the capabilities of GPT-4-turbo.
7 Conclusion
This study significantly advances the field of cybersecurity by providing a robust methodology for detecting covert communications within IPv6 networks. By leveraging a Python-based tool inspired by pcapstego, we curated a balanced dataset aimed at machine learning applications that included both normal and covertly modified IPv6 traffic. Our comparative analysis across various machine learning models—including RF, XGB, LGB, GB, DNN, CNN, RNN, GCN, and LSTM—reveals that Random Forest (RF) exhibited the most optimal performance, whereas the Graph Convolutional Network (GCN) showed the least favorable outcomes.
7.1 Contributions and Distinguishing Features
@@ 7.2 Future Directions
  • 4.
    Integration of Generative AI: Because generative AI generates and refines detection scripts dynamically based on changing threat landscapes, it can help with both the detection and classification phases. This robust tool can be used to counter covert communications because of its capacity to process and synthesize large amounts of packet data, identify subtle anomalies in IPv6 fields, and iteratively improve detection methods. Furthermore, the ability to select from a variety of machine learning models or extensive language models for classification allows the generative AI agent to swiftly adjust to newly discovered covert channel types. The integration of AI-driven real-time monitoring systems, in which generative AI creates and modifies detection algorithms on the fly based on real-time network traffic, may be the subject of future research. This would improve the overall resilience of cybersecurity defenses by enabling dynamic adaptation to new attack vectors in IPv6 environments. To further enhance the process of detecting and classifying covert communications, multi-agent systems—where various generative AI models cooperate to exchange insights on vulnerabilities—might be used.
In conclusion, our work not only enhances the detection and classification of covert communications within IPv6 networks but also sets a new standard for dataset realism and applicability in cybersecurity research. This foundation encourages further exploration into effective and adaptable cybersecurity measures in an ever-evolving digital landscape.
8 Acknowledgement
The authors thank the Department of Systems and Industrial Engineering at the University of Arizona for matching collaboration opportunities with the National Security Agency through the SIE472/572 Information Security and Research and INSuRE programs.

### v2

@@ VI-A Binary Classification
In contrast, the Graph Convolutional Network (GCN) exhibited lower performance across all evaluation metrics, suggesting limited effectiveness in the binary classification setting. This outcome may be attributed to the model’s dependence on non-Euclidean data structures, which are less prominent in the current task. Similarly, the Bidirectional Long Short-Term Memory (BiLSTM) model yielded lower scores relative to other neural network architectures, indicating that its performance in this context may be constrained and could benefit from further optimization.
VI-B Multiclass Classification
In the multiclass classification task, which presents a higher level of complexity, both the Random Forest (RF) and Extreme Gradient Boosting (XGB) models maintained relatively strong performance, suggesting their adaptability to multi-class scenarios. The Light Gradient Boosting (LGB) model achieved the highest accuracy and $F_{1}$ scores among all evaluated models, indicating its effectiveness in addressing the increased complexity of multiclass classification.
@@ VI-B Multiclass Classification
Compared the multiclass classification performance with the binary classification task, the Bidirectional LSTM (BiLSTM) demonstrated improved results in the multiclass classification setting. This relative improvement may be associated with the model’s ability to capture sequential dependencies, which can be beneficial in distinguishing between covert communication types that involve ordered patterns. For instance, in cases such as Hop Limit-based encoding, the BiLSTM appears to be effective in modeling the temporal variations in hop limit values, contributing to more accurate class differentiation. These findings suggest that sequence modeling architectures like BiLSTM may offer advantages in tasks where temporal or structural dependencies are relevant.
@@
TABLE III: Model Performance: Accuracy and $F_{1}$ Score for Binary and Multiclass Classification
However, the GCN showed a further decrease in performance in the multiclass classification setting, suggesting limited applicability in this context. This may be due to the model design, which is optimized for graph-structured data rather than the tabular feature-based representations used in covert communication detection tasks. In addition, DNN, CNN, and RNN exhibited stable and comparable performance, with CNN achieving slightly better results among neural network-based models. Although these architectures performed competitively, the improved performance of the Bi-LSTM highlights its relative strength in capturing sequential dependencies that are relevant to distinguish between different types of covert communication.
In general, traditional tree-based models, such as RF, XGB, and especially LGB, demonstrated exceptional performance in both binary and multiclass classification tasks, reaffirming their robustness for detection of covert communication. Neural network-based models also performed competitively, though GCN’s architectural design constraints led to suboptimal results. Table III summarizes the accuracy and $F_{1}$ scores for both tasks.
VII Conclusion
This study presents a systematic analysis of the potential use of various IPv6 header fields for covert communication. It also introduces machine learning-based approaches for binary detection and multiclass classification of covert traffic within IPv6 networks. Using a Python-based tool inspired by pcapstego, we curated a balanced data set aimed at machine learning applications that included normal and covertly modified IPv6 traffic. Our comparative analysis across various machine learning models, including RF, XGB, LGB, GB, DNN, CNN, RNN, GCN, and LSTM, reveals that Random Forest (RF) exhibited the most optimal performance, whereas the Graph Convolutional Network (GCN) showed the least favorable outcomes.
VII-A Contributions and Distinguishing Features
@@ VII-B Future Directions
  • Integration of Generative AI: Because generative AI generates and refines detection scripts dynamically based on changing threat landscapes, it can help with both the detection and classification phases. This robust tool can be used to counter covert communications because of its capacity to process and synthesize large amounts of packet data, identify subtle anomalies in IPv6 fields, and iteratively improve detection methods. Furthermore, the ability to select from a variety of machine learning models or extensive language models for classification allows the generative AI agent to swiftly adjust to newly discovered covert channel types. The integration of AI-driven real-time monitoring systems, in which generative AI creates and modifies detection algorithms on the fly based on real-time network traffic, may be the subject of future research. This would improve the overall resilience of cybersecurity defenses by enabling dynamic adaptation to new attack vectors in IPv6 environments. To further enhance the process of detecting and classifying covert communications, multi-agent systems that leverage the collaboration of various generative AI models to exchange insights on vulnerabilities may be employed.
In conclusion, this work improves the detection and classification of covert communications within IPv6 networks by combining protocol field analysis with machine learning techniques. The data set and methodology used aim to reflect realistic scenarios, supporting future research on practical and adaptable cybersecurity approaches.
Declarations
This study makes use of the 2019 CAIDA IPv6 Launch Day Anonymized Internet Traces for normal network traffic. Access to these traces is subject to CAIDA’s usage agreement and is not publicly available; interested researchers may request access via the CAIDA website (https://www.caida.org).
To support the reproducibility of our experiments, we have made representative samples publicly available in this GitHub repository, including both normal traffic and covert traffic samples generated during the study.
Acknowledgment
This work is supported by the National Science Foundation (NSF) under project numbers 1624668, 1921485, and 2335046; the U.S. Department of Energy–National Nuclear Security Administration (DOE-NNSA) under Award Number DE-NA0003946; and the AGILITY project (Grant No. 4263090), sponsored by the Korea Institute for Advancement of Technology (KIAT), South Korea. The authors would like to thank the Department of Systems and Industrial Engineering at the University of Arizona for matching collaboration opportunities with the National Security Agency through the Information Security Research and Education (INSuRE) research collaborative.